11x+5y+23=0
7x-2y-0=0
please reply soon as possible

Answer by stanbon(75887)

11x+5y+23=0
7x-2y-0=0
-------------
Rewrite:
11x + 5y = -23
7x - 2y = 0
--------------
Multiply the 1st equation by 2 and the 2nd by 5 to get:
22x + 10y = -46
35x - 10y = 0
-------------------
Add the two equations and solve for "x":
57x = -46
x = -0.8070...
-----------
Substitute in one of the original equations to solve for "y".
===========================================
Cheers,
Stan H.